OpenStudy (melissa_something):

Quadratic Formula question :( Will medal <3

OpenStudy (melissa_something):

\[-2\pm \sqrt{2^2}-4(1)(-4) /2\]

OpenStudy (melissa_something):

This is so aggravating. The answer is \[-1\pm \sqrt{5}\] but idk how to get it.

OpenStudy (anonymous):

All you have todo is solve what's under the sqrt

OpenStudy (anonymous):

DO you get it ?

OpenStudy (melissa_something):

I know that, I just don't know what else to do. Im stuck at the sqrt of 20

OpenStudy (anonymous):

Ok so you know quadratic formula right?

OpenStudy (melissa_something):

Yes I already plugged it in!

OpenStudy (anonymous):

just solve it. where are you stuck?

OpenStudy (melissa_something):

\[-2\pm \sqrt{20}/2\]

OpenStudy (anonymous):